If a=2^(1/3) - 2^(1/3) then find the value of2a^3 + 6a -3
 

Grade:9

2 Answers

Arun
5 years ago
Dear Arnav
 
As you see that
a = 2^(1/3) - 2^(1/3) = 0
Hence
2a³ +6a -3 
 = 2 *0 + 6 *0 -3
 = -3

Regards
Arun (askIITians forum expert)
Arnav garg
5 years ago
But the answer is 0
(All all all for do all do all em so taking so ye we ye an ye so it do it do ye to you soon so ye we it do is to get the in I'm ok with do)
Sorry for all this because I need to type more than 100 char